# Citation: Kuric E., Demcak P., Krajcovic M. (2025) Talking Surveys: How Photorealistic Embodied Conversational Agents Shape Response Quality, Engagement, and Satisfaction

### v1

@@ 1 Introduction
To address this issue, we conduct a comprehensive study with the primary goal of comparing the results of card sorting depending on whether participants sorted a full list of cards or its subset. As a sub-issue, randomized subsets in card sorting may have negative impacts on the sample size needed to obtain reliable results. To obtain empirical findings about optimal sample size, we explore the relationship between the number of participants and sorting results. As a second sub-issue, the natural diversity in participants’ psychological profiles—personality or cognitive reflection—may introduce variability detrimental for the use of randomized subsets. Representation of specific profiles may also be beneficial, leading us to investigate the effects of personality and cognitive reflection. This is a topic yet to be examined for card sorting in general, in spite of personality and cognitive reflection having established effects on user behavior, experience and perceived usability [Hattie Wilczewski and Kortum, 2024, Liapis et al., 2019, Özbek et al., 2014, Mosleh et al., 2021], as well as relationships with mental models in other contexts [Vˆırgă et al., 2014, Fisher et al., 2012].
@@ 2.1 Big Five personality model
Big Five consists of five traits: Negative Emotionality, Extraversion, Agreeableness, Conscientiousness and Open-Mindedness, associated with the following characteristics [Zekry and McKee, 2023, Tatnall, 2020]:
@@ 3.2 Procedure
Participants completed a study in the Card Sorting tool which was thus aligned with the tools standard procedural framework (see 3.3.1 Card Sorting tool). The pre-study questionnaire contained a demographic questionnaire to monitor the sample’s distribution and a Cognitive Reflection Test (CRT) to measure cognitive capacity. The post-study questionnaire contained the personality assessment questionnaire (BFI-2-XS) [Soto and John, 2017a]. Both questionnaires also incorporated an attention check question to verify participants’ engagement and effort during the study. This helped us identify participants who may exhibit straight-lining behavior in the online environment [Krajcovic et al., 2025].
@@ 3.3.3 Big Five Inventory-2
To minimize potential effects of questionnaires on engagement and thus on user behavior during card sorting, we used the extra short form of the Big Five Inventory-2 (BFI-2-XS) [Soto and John, 2017a] to assess Big Five personality traits. While the standard BFI-2 is a robust instrument for measuring the five personality traits and their respective three facets [Soto and John, 2017b], it contains 60 Likert scale items, which is considerable when the efficiency of data collection is a concern. Card sorting itself is already an effort-intensive process. As such, the 15-item BFI-2-XS was chosen to prevent fatigue from participants. In spite of its short length, BFI-2-XS captures the Big Five traits—Extraversion, Conscientiousness, Open-Mindedness, Agreeableness and Negative Emotionality—with 80% of the variability of BFI-2 [Soto and John, 2017a].
@@ 4 Results
Across both card sets, a higher number of cards elevated the time spent on card sorting. For the e-commerce card set, the difference was only moderate at median 633 seconds ($IQR:477\textendash 778$) in the Full-E group and 543 seconds ($IQR:404\textendash 783$) in the Subset-E group ($U(80)=953.5,z=1.48,p=.14$). The difference was more significant for the banking set at 748.5 seconds ($IQR:584.75\textendash 1324.25$) in the Full-B group and 539.5 seconds ($IQR:457.75\textendash 757$) in the Subset-B group ($U(80)=1126.5,z=3.14,p=.001,r=.35$).
A similar pattern was observed in the number of created categories. Within both groups Full-E and Subset-E recorded a consistent median at 5.5 ($IQR:4\textendash 7$) and no significant difference ($z=.43,p=.66$). However, between groups Full-B at 7.5 ($IQR:6\textendash 9$) and Subset-B at 6 ($IQR:4\textendash 7$), the number of created categories was significantly different ($U(80)=1107,z=2.95,p=.003,r=.33$).

### v2

# Card Sorting with Fewer Cards and the Same Mental Models? A Re-examination of an Established Practice

@@ 1 Introduction
To address this issue, we conduct a comprehensive study with the primary goal of comparing the results of card sorting depending on whether participants sorted a full list of cards or its subset. As a sub-issue, randomized subsets in card sorting may have negative impacts on the sample size needed to obtain reliable results. To obtain empirical findings about optimal sample size, we explore the relationship between the number of participants and sorting results. As a second sub-issue, the natural diversity in participants’ psychological profiles—personality or cognitive reflection—may introduce variability detrimental for the use of randomized subsets. Representation of specific profiles may also be beneficial, leading us to investigate the effects of personality and cognitive reflection. This is a topic yet to be examined for card sorting in general, in spite of personality and cognitive reflection having established effects on user behavior, experience and perceived usability [Hattie Wilczewski and Kortum, 2024, Liapis et al., 2019, Özbek et al., 2014, Mosleh et al., 2021], as well as relationships with mental models in other contexts [Vîrgă et al., 2014, Fisher et al., 2012].
@@ 2.1 Big Five personality model
Big Five consists of five traits: Negative Emotionality, Extraversion, Agreeableness, Conscientiousness and Open-Mindedness, associated with the following characteristics [D. A. Zekry and G. T. McKee (2023), 7]:
@@ 3.2 Procedure
Participants completed a study in the Card Sorting tool which was thus aligned with the tools standard procedural framework (see 3.3.1 Card Sorting tool). The pre-study questionnaire contained a demographic questionnaire to monitor the sample’s distribution and a Cognitive Reflection Test (CRT) to measure cognitive capacity. The post-study questionnaire contained the personality assessment questionnaire (BFI-2-XS) [Soto and John, 2017b]. Both questionnaires also incorporated an attention check question to verify participants’ engagement and effort during the study. This helped us identify participants who may exhibit straight-lining behavior in the online environment [Krajcovic et al., 2025].
@@ 3.3.3 Big Five Inventory-2
To minimize potential effects of questionnaires on engagement and thus on user behavior during card sorting, we used the extra short form of the Big Five Inventory-2 (BFI-2-XS) [Soto and John, 2017b] to assess Big Five personality traits. While the standard BFI-2 is a robust instrument for measuring the five personality traits and their respective three facets [Soto and John, 2017a], it contains 60 Likert scale items, which is considerable when the efficiency of data collection is a concern. Card sorting itself is already an effort-intensive process. As such, the 15-item BFI-2-XS was chosen to prevent fatigue from participants. In spite of its short length, BFI-2-XS captures the Big Five traits—Extraversion, Conscientiousness, Open-Mindedness, Agreeableness and Negative Emotionality—with 80% of the variability of BFI-2 [Soto and John, 2017b].
@@ 4 Results
Across both card sets, a higher number of cards elevated the time spent on card sorting. For the e-commerce card set, the difference was only moderate at median 633 seconds (IQR:477778IQR:477–778) in the Full-E group and 543 seconds (IQR:404783IQR:404–783) in the Subset-E group ($U(80)=953.5,z=1.48,p=.14$). The difference was more significant for the banking set at 748.5 seconds (IQR:584.751324.25IQR:584.75–1324.25) in the Full-B group and 539.5 seconds (IQR:457.75757IQR:457.75–757) in the Subset-B group ($U(80)=1126.5,z=3.14,p=.001,r=.35$).
A similar pattern was observed in the number of created categories. Within both groups Full-E and Subset-E recorded a consistent median at 5.5 (IQR:47IQR:4–7) and no significant difference ($z=.43,p=.66$). However, between groups Full-B at 7.5 (IQR:69IQR:6–9) and Subset-B at 6 (IQR:47IQR:4–7), the number of created categories was significantly different ($U(80)=1107,z=2.95,p=.003,r=.33$).